# FlightBERT++: A Non-autoregressive Multi-Horizon Flight Trajectory Prediction Framework

### v2

@@ Abstract
Flight Trajectory Prediction (FTP) is an essential task in Air Traffic Control (ATC), which can assist air traffic controllers in managing airspace more safely and efficiently. Existing approaches generally perform multi-horizon FTP tasks in an autoregressive manner, thereby suffering from error accumulation and low-efficiency problems. In this paper, a novel framework, called FlightBERT++, is proposed to i) forecast multi-horizon flight trajectories directly in a non-autoregressive way, and ii) improve the limitation of the binary encoding (BE) representation in the FlightBERT. Specifically, the FlightBERT++ is implemented by a generalized encoder-decoder architecture, in which the encoder learns the temporal-spatial patterns from historical observations and the decoder predicts the flight status for the future horizons. Compared with conventional architecture, an innovative horizon-aware contexts generator is dedicatedly designed to consider the prior horizon information, which further enables non-autoregressive multi-horizon prediction. Moreover, a differential prompted decoder is proposed to enhance the capability of the differential predictions by leveraging the stationarity of the differential sequence. The experimental results on a real-world dataset demonstrated that the FlightBERT++ outperformed the competitive baselines in both FTP performance and computational efficiency.
Introduction
@@
Flight Trajectory Prediction (FTP) is essential for Air Traffic Management (ATM), enabling many critical applications to help Air Traffic Controllers (ATCOs) manage airspace more safely and efficiently, such as traffic flow prediction (Lin, wei Zhang, and Liu 2019; Yan et al. 2023), conflict detection (Liu and Hwang 2011; Chen, Guo, and Lin 2020; Li et al. 2019), and arrival time estimation (Wang, Liang, and Delahaye 2018; Ma et al. 2023). In this context, the FTP technique has gathered more attention from researchers and achieved significant progress in the past decade.
In general, FTP in the Air Traffic Control (ATC) domain aims to forecast the flight status in future time steps according to the observed historical flight trajectories. Currently, the short FTP tasks are commonly formulated as a time series forecasting problem, which predicts future flight trajectories based on real-time observations. According to modeling approaches, the FTP methods can be classified into three distinct categories: physical models, filter-based models, and data-driven models. Physical models typically establish a set of mathematical equations based on kinematics and aerodynamics assumptions to estimate the future flight status (Tang et al. 2015; Benavides et al. 2014; FAA 2010), while the filter-based models estimate the flight trajectory iteratively via a predefined system model by considering real-time measurements (Kalman 1960; Yan et al. 2013; Thipphavong et al. 2013). However, limited by the fixed inference rules, these approaches easily suffer from error accumulation problems and not be suitable for multi-horizon FTP tasks. In contrast, data-driven models aim to learn flight transition patterns from historical data, which have emerged as the predominant approach in modern ATC applications due to remarkable performance (Wang, Liang, and Delahaye 2017; Zeng et al. 2020).
@@
Among the data-driven models, the FlightBERT is a pioneering effort that proposes binary encoding (BE) representation to encode the attributes of the trajectory points into a set of binary vectors, further formulating the FTP task as multiple binary classification (MBC) problem (Guo et al. 2023). Benefiting from the innovative idea of BE representation, the FlightBERT framework not only enhances the semantic representation of the trajectory points, but also avoids the vulnerability impacted by the normalization algorithms. However, two primary limitations should be further addressed to enhance the performance of the FTP task.
@@
  • A limitation of the BE representation is that the high-bit prediction errors will lead to outliers in the predictions. For instance, given the BE representation "0110 0100" (decimal 100), the absolute error is 128 (decimal) if the misclassification occurs in the $8^{th}$ bit ("1110 0100") while that is 1 (decimal) if the prediction error occurs in $1^{st}$ bit ("0110 0101").
@@
  • The FlightBERT performs multi-horizon prediction recursively, i.e., predicts the flight status of the next time step based on observation and iteratively applies the predicted values as pseudo-observation to obtain multi-horizon predictions. It is easy to suffer from larger accumulative errors since the prediction errors in the pseudo-observation will be accumulated during the recursive inference process. Additionally, the computational efficiency is limited by the prediction horizon due to the step-by-step prediction paradigm. As the prediction horizon increases, the inference speed will experience a severe reduction, leading to unacceptable delays for real-time applications (e.g., conflict detection).
@@
Focusing on the aforementioned challenges, in this paper, a non-autoregressive multi-horizon FTP framework, named FlightBERT++, is proposed to improve the performance of the FTP task. Thanks to the superior trajectory representation ability of the FlightBERT, the proposed framework inherits the BE representation from the FlightBERT, and is also implemented based on the MBC paradigm.
@@
In order to overcome the outlier predictions resulting from the high-bit prediction errors in the BE representation, a differential prediction paradigm is introduced into the FlightBERT++ framework. Specifically, instead of predicting the absolute values, the differential values of the trajectory attributes are formulated as the output objective in the proposed framework, which can be encoded into BE representations in less bits. In this way, the FlightBERT++ framework is able to effectively mitigate the occurrence of extremely unreliable outliers in the predictions.
@@
To achieve high-accuracy and -efficiency multi-horizon FTP prediction, a novel sequence-to-sequence (Seq2Seq) based encoder-decoder architecture is proposed to implement the FlightBERT++ framework. Specifically, the FlightBERT++ is constructed by a cascading trajectory encoder, Horizon-aware Context Generator (HACG), and Differential-prompted Decoder (DPD). The trajectory encoder encodes the observations into a trajectory-level representation, while the HACG is designed to produce multi-horizon context representations of the predicted horizons. Benefiting from the proposed HACG module, unlike conventional Seq2Seq architecture, the proposed framework can generate multi-horizon predictions directly (non-autoregressive) rather than perform recursive inference. In succession, the differential-prompted decoder is innovatively proposed to perform the high-confidence predictions based on the multi-horizon context representations, which employ the differential sequence of the observations as the prompting. The Transformer blocks are applied to build the backbone network for both the trajectory encoder and DPD, enabling non-autoregressive inference in the temporal modeling process. In this way, the proposed framework can not only mitigate the error accumulation but also yield a substantial improvement in computational efficiency for multi-horizon FTP tasks.
@@
The proposed framework is evaluated on a real-world flight trajectory dataset from an industrial ATC system. To validate the effectiveness and efficiency of the proposed framework, several competitive baselines are selected to conduct comprehensive comparisons. In addition, extensive ablation studies and insightful analysis are also performed to confirm all proposed technical improvements. The experimental results consistently demonstrated that the proposed framework efficiently addresses the outliers and error accumulation problems, outperforming the comparative baselines in both FTP accuracy and efficiency. In summary, the contributions and novelty of this work are listed as follows:
@@
  • The differential prediction paradigm is dedicatedly designed to overcome the limitations of the high-bit prediction errors in the BE representation.
@@
  • A HACG is innovatively proposed to generate multi-horizon context representations by leveraging prior horizon knowledge, which is the key to supporting non-autoregressive predictions.
@@
  • Considering the stationarity of the differential sequence in flight trajectory, a differential-prompted decoder is proposed to facilitate the learning of transition patterns in trajectory sequences, which further improves the performance of the FlightBERT++.
Methodology
@@
Problem Formulation
In general, the short-term FTP task can be formulated as a multi-variable time-series forecasting problem, i.e., predicting the aircraft status of several future horizons based on historical observations. Let an observed trajectory point $p_{t}$ in timestamp $t$, the multi-horizon FTP aims to forecast the $P_{t+1:t+n}=\{p_{t+1},p_{t+2},...,p_{t+n}\}$ based on the observation sequence $O_{t-k+1:t}=\{p_{t-k+1},...,p_{t-1},p_{t}\}$, as Eq. (1):
$P_{t+1:t+n}=\{p_{t+1},p_{t+2},...,p_{t+n}\}=\mathcal{F}(O_{t-k+1:t})$ (1)
where the $n,k$ are the number of prediction horizons and the observed sequence length, respectively. $\mathcal{F}(\cdot)$ denotes the learnable FTP model. Compared to the conventional iterative multi-horizon FTP approaches, $\mathcal{F}(\cdot)$ outputs the multi-horizon predictions directly in the FlightBERT++.
In this work, the trajectory point $p_{t}$ is formulated as a collection of six key attributes for aircraft status, as Eq. (2):
$p_{t}=[Lon_{t},Lat_{t},Alt_{t},{Vx}_{t},{Vy}_{t},{Vz}_{t}]$ (2)
where the $Lon_{t},Lat_{t},Alt_{t},{Vx}_{t},{Vy}_{t},{Vz}_{t}$ represents the longitude, latitude, altitude, and velocity in $x,y,z$ dimensions (corresponds to the longitudinal, latitudinal, and altitudinal attributes), respectively.
The Proposed Framework
Overview
The proposed framework is illustrated in Figure 1. The neural architecture of the proposed framework is cascaded by trajectory encoder, horizon-aware context generator, and differential-prompted decoder. Given the observed trajectory sequence $O_{t-k+1:t}$, the object of the trajectory encoder is to learn the temporal-spatial correlations of the observations and abstract them to a high-dimensional representation $\mathbf{Traj}_{enc}$, as Eq. (3).

### v4

# A Non-autoregressive Multi-Horizon Flight Trajectory Prediction Framework with Gray Code Representation

@@ Abstract
Flight Trajectory Prediction (FTP) is an essential task in Air Traffic Control (ATC), which can assist air traffic controllers in managing airspace more safely and efficiently. Existing approaches generally perform multi-horizon FTP tasks in an autoregressive manner, thereby suffering from error accumulation and low-efficiency problems. In this paper, a novel framework, called FlightBERT++, is proposed to i) forecast multi-horizon flight trajectories directly in a non-autoregressive way, and ii) improve the limitation of the binary encoding (BE) representation in the FlightBERT framework. Specifically, the proposed framework is implemented by a generalized encoder-decoder architecture, in which the encoder learns the temporal-spatial patterns from historical observations and the decoder predicts the flight status for the future horizons. Compared to conventional architecture, an innovative horizon-aware contexts generator is dedicatedly designed to consider the prior horizon information, which further enables non-autoregressive multi-horizon prediction. Additionally, the Gray code representation and the differential prediction paradigm are designed to cope with the high-bit misclassifications of the BE representation, which significantly reduces the outliers in the predictions. Moreover, a differential prompted decoder is proposed to enhance the capability of the differential predictions by leveraging the stationarity of the differential sequence. Extensive experiments are conducted to validate the proposed framework on a real-world flight trajectory dataset. The experimental results demonstrated that the proposed framework outperformed the competitive baselines in both FTP performance and computational efficiency. The code is publicly available at: https://github.com/gdy-scu/FlightBERT_PP_V2.
Index Terms:
Flight trajectory prediction, Gray code representation, Horizon-aware contexts generator, Multi-horizon forecasting, Non-autoregressive.
I Introduction
@@
Flight Trajectory Prediction (FTP) is essential for Air Traffic Management (ATM), enabling many critical applications to help Air Traffic Controllers (ATCOs) manage airspace more safely and efficiently, such as traffic flow prediction [1, 2], conflict detection [3, 4, 5], and arrival time estimation [6, 7]. In particular, the FTP is one of the fundamental techniques to support the Trajectory-Based Operation (TBO) that is being promoted by both the Single European Sky ATM Research (SESAR) and the New Generation Air Transportation Systems (NextGen) [8]. In this context, the FTP technique has gathered more attention from researchers and achieved significant progress in the past decade.
@@
In general, FTP in the Air Traffic Control (ATC) domain aims to forecast the flight status in future time steps according to the observed historical flight trajectories. Accurate FTP requires high prediction capacity to capture the temporal-spatial dependencies between the observations and predictions. To investigate the robust representations for the trajectory attributes, the binary encoding (BE) is proposed to encode the attributes of the trajectory points into a set of binary vectors in our previous FlightBERT framework [9], further formulating the TP task as multiple binary classification (MBC) problem. Benefiting from the innovative idea and dedicatedly network design, the FlightBERT framework not only enhances the semantic representation of the trajectory points, but also avoids the vulnerability impacted by the normalization algorithms. However, two primary limitations should be further addressed to enhance the performance of the FTP task.
@@
  • A limitation of the Binary Encoding (BE) representation is that the high-bit misclassification will lead to outliers in the predictions [9]. For instance, given the BE representation ”0110 0100” (decimal 100), the absolute error is 128 (decimal) if the prediction error occurs in the $8^{th}$ bit (”1110 0100”) while that is 1 (decimal) if the prediction error occurs in $1^{st}$ bit (”0110 0101”).
@@
  • The FlightBERT performs multi-horizon prediction recursively, i.e., predicts the flight status of the next time step based on observation and iteratively applies the predicted values as pseudo-observation to obtain multi-horizon predictions. On the one hand, it is easy to suffer from larger accumulative errors since the prediction errors in the pseudo-observation will be accumulated during the recursive inference process. On the other hand, the computational efficiency is limited by the prediction horizon due to the step-by-step prediction paradigm. As the prediction horizon increases, the inference speed will experience a severe reduction, leading to unacceptable delays for real-time applications (e.g., conflict detection).
@@
Focusing on the aforementioned challenges, in this paper, a non-autoregressive multi-horizon flight trajectory prediction framework, named FlightBERT++, is proposed to improve the performance of the FTP task. Thanks to the superior trajectory representation ability of the BE, the proposed FlightBERT++ framework inherits the powerful representation capacity from the FlightBERT, and is also implemented based on the MBC paradigm.
@@
In order to overcome the outlier predictions resulting from the high-bit misclassification in the BE representation, in this work, an innovative Gray Code (GC) representation, and a differential prediction paradigm are incorporated into the FlightBERT++ framework. Specifically, i) the Gray code is employed to alternate the BE representation to ensure steady bit transition patterns of the trajectory attributes between the adjacent time steps. ii) instead of predicting the original values, the differential values of the trajectory attributes are formulated as the output objective in the proposed framework, which can be encoded into GC representations in fewer bits. In this way, the FlightBERT++ framework is able to effectively mitigate the occurrence of extremely unreliable outliers in the predictions.
@@
To achieve high-accuracy and -efficiency multi-horizon FTP prediction, a novel sequence-to-sequence (Seq2Seq) based encoder-decoder architecture is proposed to implement the FlightBERT++ framework. Specifically, the proposed framework is constructed by cascading the trajectory encoder, Horizon-aware Context Generator (HACG), and Differential-prompted Decoder (DPD). Firstly, the trajectory encoder encodes the observations into a trajectory-level representation, while the HACG is designed to produce multi-horizon context representations of the predicted horizons. Benefiting from the proposed HACG module, unlike conventional Seq2Seq architecture, the proposed framework can generate multi-horizon predictions directly (non-autoregressive) rather than perform recursive inference. In succession, the differential-prompted decoder is innovatively proposed to perform the high-confidence predictions based on the multi-horizon context representations, which employ the differential sequence of the observations as the prompting. In the proposed framework, the Transformer blocks are applied to build the backbone network for both the trajectory encoder and DPD, which allows the network to perform non-autoregressive inference in the temporal modeling process. In this way, the proposed framework can not only mitigate the error accumulation but also yield a substantial improvement in computational efficiency for multi-horizon FTP tasks.
This paper is an extended version of our previous work published at the AAAI Conference [10], which improves our original work as follows: i) An innovative Gray Code representation is proposed to encode the trajectory attributes to tackle the high-bit misclassification problems, which achieves 29.0% and 21.8% mean deviation error reduction in 1 and 15 horizon predictions. ii) Two more competitive baselines are considered to validate the proposed approach, including the latest work WTFTP [11] and the typical multi-horizon forecast architecture Transformer-Seq2Seq [12]. iii) Additional technical details, neural architecture and related works are provided to enhance the systematicity and reproducibility of this work. iv) More insightful experiments (as well as ablation studies), results, visualization and analyses for different flight phases are also provided in this version.
@@
The proposed framework is evaluated on a real-world flight trajectory dataset from an industrial ATC system. To validate the effectiveness and efficiency of the proposed framework, a total of 7 competitive baselines are selected to conduct comprehensive comparisons. In addition, extensive ablation studies and insightful analysis are also performed to confirm all proposed technical improvements. The experimental results consistently demonstrated that the proposed framework efficiently addresses the outliers and error accumulation problems, outperforming the baselines in both FTP accuracy and efficiency. In summary, the contributions and novelty of this work are listed as follows:
@@
  • The Gray code representation and differential prediction paradigm is dedicatedly designed to overcome the limitations of the high-bit misclassification in BE representation.
@@
  • A HACG is innovatively proposed to generate multi-horizon context representations by leveraging prior horizon knowledge, which is the key idea to supporting non-autoregressive predictions.
@@
  • Considering the stationarity of the differential sequence in flight trajectory, a differential-prompted decoder is proposed to assist in learning the transition patterns of the trajectory sequence, which further improves the performance of the FlightBERT++.
  • We conducted extensive comparison experiments on the real-world trajectory dataset to demonstrate the effectiveness of the proposed frameworks. In addition, some diagnostic studies and analyses are applied to explore the robustness and superiority of the proposed framework.
The rest of this paper is organized as follows. The related literature and works of the FTP task are briefly reviewed in Section II. Section III detailed presents the methodologies of the proposed FlightBERT++ framework. The dataset, experimental settings, baselines, and evaluation metrics are described in Section IV. The experimental results, visualization, and discussions are presented in Section V. Finally, the work of this paper is concluded in Section VI.
II Related Work
II-A Flight Trajectory Prediction
Currently, the short FTP tasks are commonly formulated as a time series forecasting (TSF) problem, which predicts future flight trajectories based on historical observations. According to modeling approaches, the FTP methods can be classified into three distinct categories: physical models, filter-based models, and data-driven models.
II-A1 Physical models
Physical models typically establish a set of mathematical equations based on kinematics and aerodynamics assumptions to estimate the future flight status [13, 14, 15, 8]. In this paradigm, the aircraft is generally modeled as a point mass and many factors are considered and incorporated into the physical models to perform more accurate predictions, such as aircraft dynamics, aircraft performance, and atmospheric conditions [16, 17, 18]. Although modern physical models can obtain the Aircraft Performance Models (APMs) and Operating Procedure Coefficients (OPCs) from the public database, such as Base of Aircraft Data (BADA) [19], certain sensitive information is still not disclosed to researchers publicly. In [20], an aircraft takeoff mass inferring method was proposed to enhance the APMs by studying the kinetic model at the lift-off moment and observing the aircraft motion on the runway. To improve the trajectory prediction accuracy, an unknown point-mass model parameters learning approach was proposed to estimate the mass and thrust of the aircraft [21]. These unknown parameters were learned from historical observations and adjusted by fitting the modeled specific power to the observed energy rate.
It is evident that physical models have been extensively investigated over the past few decades and derived various applications in modern ATC and avionics systems. However, the performance of physical models is highly sensitive to the specific performance parameters of the aircraft and typically requires modeling tailored to different aircraft types. In addition, limited by the fixed inference rules, physical models usually suffer from performance degradation during flight maneuvering phases due to more complex transition patterns.
II-A2 Filter-based models
Filter-based models are the classical FTP approaches that estimate the flight trajectory iteratively via a predefined system model by considering real-time measurements. These approaches are generally implemented based on the Kalman Filter (KF) [22] or its variants, which continuously corrects the trajectory estimates based on the observed measurements, providing a more accurate and dynamically updated prediction [23, 24]. In practice, filter-based models are commonly used for short-term FTP tasks or target tracking tasks to mitigate errors or uncertainties in the measurements. To improve the FTP performance in conflict detection and resolution applications, the KF algorithm was employed to provide the aircraft state estimation in [25]. The estimated flight state was further used in flight intent determination and predicted the future trajectory. In [26], the Unscented Kalman Filter (UKF) algorithm was applied to simulate and predict target trajectory and achieve the flying target tracking. Towards the challenges of target motion-mode uncertainty and nonlinearity, the Multiple Model (MM) methods have been generally considered the mainstream approach in the field of target tracking [27, 28]. Moreover, the hybrid estimation algorithm was also developed to track aircraft in terminal airspace [29].
Compared to physical models, since the dynamically updated nature of parameters in the system model, Filter-based models can adapt to various flight scenarios flexibly. However, these approaches are not suitable for multi-step prediction tasks as they rely on real-time observations to update the parameters of the system model.
II-A3 Data-driven models
Data-driven models aim to learn flight transition patterns from historical data, enabling high-precision predictions. These approaches usually build FTP models using machine learning or deep learning techniques and fit the model parameters in the training process [30, 31]. In [32], a stochastic FTP method was proposed to align trajectories into a 3D grid network and utilize the Hidden Markov Model (HMM) to account for environmental uncertainties. In [33], a machine learning approach was proposed to perform FTP tasks using the stepwise regression algorithm for arrival time prediction. Furthermore, the Gaussian Mixture Model (GMM) [34], Gaussian Processes [35], Support Vector Machine (SVM) [36], and various variant algorithms were also introduced and adapted into FTP tasks.
With the development of the deep neural networks (DNNs), various sequential modeling neural architectures were employed to implement the FTP tasks, such as Recurrent Neural Networks (RNNs), Transformer [12], and Convolutional Networks (CNN). In [37], the backpropagation (BP) neural network was employed to perform the high-precision four-dimensional TP task. The Long Short Term Memory (LSTM) network was also widely introduced into FTP tasks due to its significant temporal modeling capabilities [38]. A hybrid architecture was proposed to capture the spatial-temporal features of the trajectory data by combining CNN and LSTM network [39]. Moreover, the prior physics knowledge was also incorporated into deep learning based FTP approaches to enhance the model performance in different flight phases [40, 41]. In addition, the data mining and Bayesian neural network were also introduced to model FTP tasks in some ATC applications [42, 43, 44].
In general, deep learning-based models have emerged as the predominant approach in modern FTP applications owing to their remarkable performance. Nevertheless, it is worth noting that data-driven models necessitate high-quality and representative training data to achieve desired performance.
II-A4 Multi-horizon Forecasting
Currently, most existing short-term FTP works focused on capturing the flight transition patterns and performing iteratively multi-horizon inference using predicted results in the last time-step. However, the iteratively multi-horizon forecasting paradigm suffered from error accumulation problems due to the pseudo-observations as inputs. Towards this gap, the Seq2Seq architectures were applied to develop the multi-horizon FTP tasks, which are able to generate the trajectory of multiple time steps directly. In [45], an encoder-decoder architecture implemented on CNN and Gated Recurrent Units (GRU) was proposed to conduct FTP models, which exhibit a remarkable performance in multi-horizon predictions. Despite the partial alleviation of error accumulation by Seq2Seq models, the performance and inference speed remain limited in multi-horizon predictions due to the autoregressive inference principles. Therefore, this work aims to not only improve the limitations of the BE representations but also develop a non-autoregressive framework to perform high-precision multi-horizon FTP tasks.
III Methodology
III-A Problem Formulation
In general, the short-term FTP task can be formulated as a multi-variable time-series forecasting problem, i.e., predicting the trajectory point (composed of several trajectory attributes) of future horizons based on historical observations. Let an observed trajectory point $p_{t}$ in timestamp $t$, the multi-horizon FTP aims to forecast the $P_{t+1:t+n}=\{p_{t+1},p_{t+2},...,p_{t+n}\}$ based on the observation sequence $O_{t-k+1:t}=\{p_{t-k+1},...,p_{t-1},p_{t}\}$. Mathematically, it can be abstracted as Eq. (1):
$P_{t+1:t+n}=\{p_{t+1},p_{t+2},...,p_{t+n}\}=\mathcal{F}(O_{t-k+1:t})$ (1)
where the $n,k$ are the number of prediction horizons and the observed sequence length, respectively. $\mathcal{F}(\cdot)$ denotes the learnable FTP model. Compared to the conventional approaches, instead of performing the prediction iteratively, $\mathcal{F}(\cdot)$ outputs the multi-horizon predictions directly in the proposed FlightBERT++ framework.
In this work, the trajectory point $p_{t}$ is formulated as a collection of six key attributes that mainly depict the aircraft status. The definition is presented as follows.
$p_{t}=[Lon_{t},Lat_{t},Alt_{t},{Vx}_{t},{Vy}_{t},{Vz}_{t}]$ (2)
where the $Lon_{t},Lat_{t},Alt_{t},{Vx}_{t},{Vy}_{t},{Vz}_{t}$ represents the longitude, latitude, altitude, and velocity in $x,y,z$ dimensions of the trajectory point, respectively. The $x,y,z$ dimensions of the velocity correspond to the longitudinal, latitudinal, and altitudinal attributes, respectively.
III-B Gray Code Representation and Differential Prediction
In our previous work, i.e., FlightBERT [9], the binary encoding representation was proposed to convert the scalar trajectory attributes into high-dimensional $n\mbox{-}hot$ vectors to boost the FTP tasks. Consequently, FlightBERT formulated the FTP as the multi-binary classification (MBC) task. As mentioned in Section I, although FlightBERT demonstrated great performance improvements compared to conventional FTP approaches, the susceptibility of the BE representation to high-bit misclassifications inevitably leads to some outliers in the prediction results.
In practice, the flight transition patterns of civil aviation aircraft exhibit prominent stability in terms of the longitude and latitude measurements, and trend variations in trajectory attributes typically fluctuate in a small range. In the majority of instances involving the BE representation, even marginal alterations (e.g., a decimal increment of 1) in trajectory attributes in the next time step yield significant disparities between resulting $n\mbox{-}hot$ vectors, such as decimal 127 (BE: 0111 $1$) and 128 (BE: 1000 $0$). It is believed that the significant changes in bit transition patterns in the BE representations bring challenges to the model to effectively capture the trajectory dynamics.
The Gray code, also known as reflected binary code, is an ordering of the binary numeral system, such that two successive values differ in only one bit (binary digit). Inspired by this, the Gray Code (GC) representation is proposed to alternate BE to encode the trajectory attributes into $n\mbox{-}hot$ vectors, which minimizes the bit variations between consecutive time steps. Specifically, let an observed altitude sequence $Alt=[780,781,....,796,805]$, the encoding process and a detailed comparison between BE and GC representation are depicted in Fig. 1. Notably, when the values increment from 779 to 780 (decimal), the GC representations exhibit only the $3^{rd}$ bit transitioning from 1 to 0, whereas a total of 3 bits have state changes in BE representations. Leveraging this inherent characteristic of GC, the model is expected to effectively capture the bit transition patterns, thereby alleviating high-bit misclassifications.
Figure 1: Comparison of the BE, GC, and Differential GC representations. The examples of bit state changes are marked by the red rectangles.
Moreover, to further mitigate the high-bits misclassifications of the BE representation, the differential prediction paradigm is introduced into the FlightBERT++ framework, i.e., the objective of the decoder is to predict the differential GC representation instead of the GC representation raw absolute values. This design has the following merits: i) Since the trend variations of differential in the observation sequence are with higher stationary compared to the raw absolute sequence, the dynamics of the trajectory sequence can be effectively captured by the FTP model. ii) By using differential values, their numerical range is smaller, allowing for the encoded GC representation with fewer bits. Consequently, even if there are misclassifications in the higher bits, the error can be limited within a certain range, thereby eliminating the outlier in predictions. Therefore, it is believed that predicting the differential values in GC representation is an effective way to improve the performance of the FTP task.
@@ III-B Gray Code Representation and Differential Prediction
III-C The Proposed Framework
The proposed framework is illustrated in Figure 2. The neural architecture of the proposed framework is cascaded by a trajectory encoder, horizon-aware context generator, and Differential-prompted Decoder (DPD). Given the observed trajectory sequence $O_{t-k+1:t}$, the object of the trajectory encoder is to learn the temporal-spatial correlations of the observations and abstract them to a high-dimensional representation $\mathbf{Traj}_{enc}$, as Eq. (3).